the product of 7 and the sum of a number and 5 result in the difference of 5 and the number.
eighteen less than twice a number is -28. Find the number.

Answer by addingup(3677)

Let the number be x:
the product of 7 and the sum of a number and 5: 7(x+5)
result: =
in the difference of 5 and the number: 5-x
:
Put it all together:
7(x+5) = 5-x using the Distributive Property of multiplication, distribute the 7 over the elements in the parenthesis:
7x+35 = 5-x subtract 35 and add x on both sides
8x = -30
x = -3.75
-----------------------------------------------
Check:
7(-3.75+5) = 5-(-3.75)
-26.25+35 = 5+3.75
8.75 = 8.75 Correct
:
John